the chemical formula of washing soda is Na2CO3

a. what is the formula mass of washing soda?

b. what is the mass of 3.5 mol of washing soda?

c. how many formula units are there in 350g of washing soda?


1
Expert's answer
2021-11-11T14:54:02-0500

Formula unit mass of Na2

​CO3

= 2×atomic mass of Na+1×atomic mass of C+3×atomic mass of O

=2×22.98+1×12.01+3×16

=105.97u

Formula unit mass of 10H2 ​O

=10×atomic mass of H2​O

=10×18.02

=180.02u

Formula unit mass of Na2​CO3​.10H2​O

=105.97+180.02 

=285.99u

b)The molar mass of a compound defines the mass of 1 mole of that particular substance and the number of grams per mole of a compound. In other words, the molar mass is the total mass of all the atoms in grams that make a mole of a particular molecule

=3.5 x 285.99

=1004.4 g

c)Since sodium carbonate contains two atoms of sodium, one atom of carbon and three atoms of oxygen.

Na : 2 x 23.0 = 46

C : 1 x 12.0 = 12

O : 3 x 16 = 48

When we add up the total values i.e., 46 + 12 + 48 = 106

Therefore, the molar mass of Na2CO3 is 106 g/mole.